OpenStudy (anonymous):

PLEASE HELP FAN AND MEDAL FOR ANSWER What is the value of w? 5(9 − 3) = w • 9 – 5 • 3

OpenStudy (anonymous):

What is the value of w? 5(9 − 3) = w x 9 – 5 x 3

OpenStudy (anonymous):

\[5(9-3)=w \times9-5\times3~~=~~5(6)=w \times9-15\]That's the start

OpenStudy (anonymous):

what's the value of w

OpenStudy (anonymous):

\[30=x \times-5\]

OpenStudy (anonymous):

i think w=5 is the right answer plug it in for w and then solve to see if its right.